# Memorization in Self-Supervised Learning Improves Downstream Generalization

### v2

@@ Abstract
Self-supervised learning (SSL) has recently received significant attention due to its ability to train high-performance encoders purely on unlabeled data—often scraped from the internet. This data can still be sensitive and empirical evidence suggests that SSL encoders memorize private information of their training data and can disclose them at inference time. Since existing theoretical definitions of memorization from supervised learning rely on labels, they do not transfer to SSL. To address this gap, we propose SSLMem, a framework for defining memorization within SSL. Our definition compares the difference in alignment of representations for data points and their augmented views returned by both encoders that were trained on these data points and encoders that were not. Through comprehensive empirical analysis on diverse encoder architectures and datasets we highlight that even though SSL relies on large datasets and strong augmentations—both known in supervised learning as regularization techniques that reduce overfitting—still significant fractions of training data points experience high memorization. Through our empirical results, we show that this memorization is essential for encoders to achieve higher generalization performance on different downstream tasks. [^1] Code for reproducing our experiments can be found at https://github.com/sprintml/MemorizationSSL.
@@ 1 Introduction
In recent years, self-supervised learning (SSL) has emerged as a new potent learning paradigm. SSL encoders can be trained without reliance on labeled data, which is often hard and expensive to obtain. Instead, SSL leverages the existence of large amounts of unlabeled data—often scraped from the internet—to obtain state-of-the-art performance in various domains, ranging from computer vision (He et al., 2022; Chen et al., 2020; Chen & He, 2021; Caron et al., 2021) to natural language processing (Devlin et al., 2018; Radford et al., ).
@@ 2 Background and Related Work
As an important property of learning algorithms and neural networks, memorization has been actively studied in supervised learning (Zhang et al., 2016; Arpit et al., 2017; Chatterjee, 2018; Carlini et al., 2019; 2021; 2022). The fundamental idea to quantify memorization relies on the impact that a single training data point has on the predictions of the resulting models with a larger impact (or increased ”hardness” of learning the data point (Arpit et al., 2017; Sadrtdinov et al., 2021)) indicating a higher level of memorization (Feldman, 2020). While memorization has been shown to be important for generalization related properties in supervised learning (Feldman, 2020; Feldman & Zhang, 2020) or the supervised downstream classifiers within transfer learning (Bansal et al., 2020), this aspect has not been studied for SSL. The only work which considers memorization in SSL proposes the concept of Déjà Vu memorization (Meehan et al., 2023) and quantifies how much SSL encoders associate specific views (for example of background crops) with the foreground objects in training images. To assess whether an encoder exhibits Déjà Vu memorization for a given training data point, the framework obtains the representation for a crop of the data point, and compares the representation with representations of labeled data points from a public dataset that has the same distribution as the encoder’s training data. If the labels within the $k$ nearest neighbors of the crop in the representation space are highly consistent, the data point is marked as memorized. Since the core property of SSL is training without labels, the biggest limitation of the Déjà Vu memorization is the assumption about access to labeled data from the same distribution as the training set of the encoder. Additionally, Déjà Vu memorization relies on a particular SSL augmentation, namely cropping. However, not all SSL methods use cropping. Finally, SSL encoders are applied to a myriad of downstream tasks other than classification (e.g., multi-label classification, segmentation, depth detection) where the concept of a single class per input does not exist—rendering this definition of memorization narrow. Our definition of memorization is based on representations, which are output by all SSL encoders, thus it is independent of particular augmentations, downstream tasks, or the availability of auxiliary information, such as class labels.
@@ 4.3 Memorized Data Points
Additionally, we analyze what types of data points are memorized. In Figure 1, we already showed visually that, similar to supervised learning, atypical examples experience a higher memorization in SSL than standard data points. Additionally, we show in Figure 9 and Table 13 in Section B.4 that SSL and supervised learning differ notably in the data points that they assign the highest memorization scores to while the SSL setups memorize in a more similar way. Especially, the SSL setups that share the same training method or the same encoder architecture are the most consistent.
@@ 5 Conclusion
Acknowledgements
We would like to acknowledge our sponsors, who support our research with financial and in-kind contributions: CIFAR through the Canada CIFAR AI Chair program, DARPA through the GARD program, Intel, Meta, NFRF through an Exploration grant, and NSERC through the Discovery Grant and COHESA Strategic Alliance. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute. The authors also gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre (JSC). We thank all anonymous reviewers for their constructive comments. Finally, we would like to thank members of CleverHans Lab and the SprintML Lab (https://sprintml.com) for their valuable feedback.
@@ B.4 Analysis of Memorized Samples
Table 13: Results of Kendall’s Tau Test. We test the consistency of the rankings statistically of 5000 candidates over all models used for evaluation in this paper. Note that the score is symmetric. We repeat the values in gray for the reader’s convenience.
@@
Figure 9: Samples with the highest memorization over different SSL frameworks and encoder architectures. We depict per setup the top 10 data points with the highest memorization scores and their ground-truth labels.
We first visually inspect the samples that experience the highest memorization over all frameworks in Figure 9. Overall, the highest memorized samples seem to be more consistent between the different SSL frameworks than between SSL and supervised learning. This holds for both the supervised models, the one that is trained 100 epochs and the one that is trained 10 epochs, and thereby matches the downstream performance of the SSL encoders (see Table 11). To quantify this visual impression, we analyze the rankings of memorization scores over the 5000 candidates for all different setups with a pairwise Kendall’s Tau test. We present the results in Table 13. The null-hypothesis of the test is an absence of association between the two rankings, which means that when we have a p-value below $0.05$, i.e., when we can reject the null-hypothesis, there is an association in the ranking. In the table, we indeed observe that the consistency between the rankings of memorization scores between different SSL frameworks is higher than the consistency between SSL and supervised models. In addition, among different SSL frameworks, the ones that share the same architecture (or training method) have a higher consistency.
@@ D.3 Intuition Behind our Memorization Score
[…]
@@ D.4 The Link between Memorization and Generalization
[…]
@@
[…]
@@
[…]
@@
[…]

### v3

@@ Abstract
Self-supervised learning (SSL) has recently received significant attention due to its ability to train high-performance encoders purely on unlabeled data—often scraped from the internet. This data can still be sensitive and empirical evidence suggests that SSL encoders memorize private information of their training data and can disclose them at inference time. Since existing theoretical definitions of memorization from supervised learning rely on labels, they do not transfer to SSL. To address this gap, we propose SSLMem, a framework for defining memorization within SSL. Our definition compares the difference in alignment of representations for data points and their augmented views returned by both encoders that were trained on these data points and encoders that were not. Through comprehensive empirical analysis on diverse encoder architectures and datasets we highlight that even though SSL relies on large datasets and strong augmentations—both known in supervised learning as regularization techniques that reduce overfitting—still significant fractions of training data points experience high memorization. Through our empirical results, we show that this memorization is essential for encoders to achieve higher generalization performance on different downstream tasks.
@@ 1 Introduction
In recent years, self-supervised learning (SSL) has emerged as a new potent learning paradigm. SSL encoders can be trained without reliance on labeled data, which is often hard and expensive to obtain. Instead, SSL leverages the existence of large amounts of unlabeled data—often scraped from the internet—to obtain state-of-the-art performance in various domains, ranging from computer vision (He et al., 2022; Chen et al., 2020; Chen & He, 2021; Caron et al., 2021) to natural language processing (Devlin et al., 2018; Radford et al., ; Brown et al., 2021).
@@ 2 Background and Related Work
As an important property of learning algorithms and neural networks, memorization has been actively studied in supervised learning (Zhang et al., 2016; Arpit et al., 2017; Chatterjee, 2018; Carlini et al., 2019; 2021; 2022). The fundamental idea to quantify memorization relies on the impact that a single training data point has on the predictions of the resulting models with a larger impact (or increased ”hardness” of learning the data point (Arpit et al., 2017; Sadrtdinov et al., 2021)) indicating a higher level of memorization (Feldman, 2020). Brown et al. (2021) demonstrate that models must encode information contained in a large number of training examples in order to achieve high accuracy. To arrive at this conclusion, they study supervised learning (next-symbol prediction and classification) whereas we study SSL. Additionally, while memorization has been shown to be important for generalization related properties in supervised learning (Feldman, 2020; Feldman & Zhang, 2020) or the supervised downstream classifiers within transfer learning (Bansal et al., 2020), this aspect has not been studied for SSL. The only work which considers memorization in SSL proposes the concept of Déjà Vu memorization (Meehan et al., 2023) and quantifies how much SSL encoders associate specific views (for example of background crops) with the foreground objects in training images. To assess whether an encoder exhibits Déjà Vu memorization for a given training data point, the framework obtains the representation for a crop of the data point, and compares the representation with representations of labeled data points from a public dataset that has the same distribution as the encoder’s training data. If the labels within the $k$ nearest neighbors of the crop in the representation space are highly consistent, the data point is marked as memorized. Since the core property of SSL is training without labels, the biggest limitation of the Déjà Vu memorization is the assumption about access to labeled data from the same distribution as the training set of the encoder. Additionally, Déjà Vu memorization relies on a particular SSL augmentation, namely cropping. However, not all SSL methods use cropping. Finally, SSL encoders are applied to a myriad of downstream tasks other than classification (e.g., multi-label classification, segmentation, depth detection) where the concept of a single class per input does not exist—rendering this definition of memorization narrow. Our definition of memorization is based on representations, which are output by all SSL encoders, thus it is independent of particular augmentations, downstream tasks, or the availability of auxiliary information, such as class labels.
@@ 4.3 Memorized Data Points
Additionally, we analyze what types of data points are memorized. In Figure 1, we already showed visually that, similar to supervised learning, atypical examples experience a higher memorization in SSL than standard data points. Additionally, we show in Figure 9 and Table 13 in Section B.4 that SSL and supervised learning differ notably in the data points that they assign the highest memorization scores to while the SSL setups memorize in a more similar way. Especially, the SSL setups that share the same training method or the same encoder architecture are most consistent.
@@ 5 Conclusion
Acknowledgments
We would like to acknowledge our sponsors, who support our research with financial and in-kind contributions: Amazon, Apple, CIFAR through the Canada CIFAR AI Chair, DARPA through the GARD project, Intel, Meta, NSERC through the Discovery Grant, the Ontario Early Researcher Award, and the Sloan Foundation. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute.
@@ B.4 Analysis of Memorized Samples
Table 13: Results of Kendall’s Tau Test. We test consistency of the rankings statistically of 5000 candidates over all models used for evaluation in this paper. Note that the score is symmetric. We repeat the values in gray for the reader’s convenience.
@@
Figure 9: Samples with highest memorization over different SSL frameworks and encoder architectures. We depict per setup the top 10 data points with the highest memorization scores and their ground-truth labels.
We first visually inspect the samples that experience the highest memorization over all frameworks in Figure 9. Overall, the highest memorized samples seem to be more consistent between the different SSL frameworks than between SSL and supervised learning. This holds for both the supervised models, the one that is trained 100 epochs and the one that is trained 10 epochs, and thereby matches downstream performance of the SSL encoders (see Table 11). To quantify this visual impression, we analyze the rankings of memorization scores over the 5000 candidates for all different setups with a pairwise Kendall’s Tau test. We present the results in Table 13. The null-hypothesis of the test is an absence of association between the two rankings, which means that when we have a p-value below $0.05$, i.e., when we can reject the null-hypothesis, there is an association in the ranking. In the table, we indeed observe that the consistency between the rankings of memorization scores between different SSL frameworks is higher than the consistency between SSL and supervised models. In addition, among different SSL frameworks, the ones that share the same architecture (or training method) have a higher consistency.
@@ D.3 Intuition Behind our Memorization Score
We say that an encoder $f$ satisfies $c$-strong alignment on datapoint $x_{i}$ if $\;\forall\;x^{\prime},x^{\prime\prime}\in\text{Aug}(x_{i})$, $d(f(x^{\prime}),f(x^{\prime\prime}))\leq c$.
@@ D.4 The Link between Memorization and Generalization
We say that the augmentation set of a datapoint $z$ satisfies $\sigma$-overlap if there exists a region $\text{Aug}^{0}(z)\subseteq\text{Aug}(z)$ which overlaps with the augmentation set of a training datapoint $x_{i}\in S$ and so that $P[b\in\text{Aug}^{0}(x)]\geq\sigma\cdot P[b\in\text{Aug}(x)]$.
@@
We say that a datapoint $z$ is $\beta$-close to a training datapoint $x_{i}\in S$ if $\min_{x_{i}^{\prime}\in\text{Aug}(x_{i})}||x_{i}^{\prime}-z||=\beta$.
@@
Given an encoder $f$ satisfying $c$-strong alignment over point $x$ and a test datapoint $z_{i}\in S_{\text{test}}$ which satisfies $\beta_{i}$-closeness to point $x$, $d(f(x),f(z_{i}))\leq L\beta_{i}+c$
@@
Given an encoder $f$ with $\mathcal{L}_{\text{align}}(f,x)\leq c$ and point $z_{i}$ which satisfies $\sigma$-overlap with point $x$, $\mathcal{L}_{\text{align}}(f,z_{i})\leq\sigma\cdot c+(1-\sigma)\cdot L\cdot\mathbb{E}_{z^{\prime},z^{\prime\prime}\in(\text{Aug}(z_{i})\setminus\text{Aug}(z_{i})\cap\text{Aug}(x))}||z^{\prime}-z^{\prime\prime}||$